Find the number of ways in which 9 people can be divided into 2 groups: the first group has 5 people and the second group has 4 people.

Answer by stanbon(75887)

Find the number of ways in which 9 people can be divided into 2 groups: the first group has 5 people and the second group has 4 people.
-----
# of 4 people groups = # of 5 people groups if the total # of people is 9
9C4 = (9*8*7*6)/(1*2*3*4) = (3*1*7*6) = 21*6 = 126
Cheers,
Stan H.
-----------